OpenStudy (emogirl100):

Identify the factors of x2 − 4x − 21. (x − 7)(x + 3) (x + 7)(x − 3) (x + 21)(x − 1) (x − 21)(x + 1)

jhonyy9 (jhonyy9):

x^2 -7x +3x -7*3 = ? do you see it now what and from where you can factorizing out ?

OpenStudy (emogirl100):

is it a

jhonyy9 (jhonyy9):

@emogirl100 the rule of OS - please - do not give never directly answers

OpenStudy (emogirl100):

then is it (x − 7)(x + 3)

OpenStudy (emogirl100):

ugh i did my own work i just want to know if i got it right jezz @jhonyy9

OpenStudy (emogirl100):

@Zale101 plz help

OpenStudy (sshayer):

correct.

OpenStudy (emogirl100):

thx you they gave me a warning because he didnt help me @sshayer

OpenStudy (emogirl100):

uuuugggggghhhhhhhhh

OpenStudy (sshayer):

can,t say anything.

jhonyy9 (jhonyy9):

sorry i was occuped with one begined help - but you are right sure

OpenStudy (emogirl100):

@jhonyy9 next time just say nay or yay because i got that questoin almost wrong

OpenStudy (emogirl100):

thx u

jhonyy9 (jhonyy9):

sorry than i understand you right so this question was wrong ? just bc. x^2 -7x +3x -7*3 = ? from thes you get x*x -7x +3x -7*3 = x(x-7) +3(x-7) = (x+3)(x-7) this is the right answer sure